There are 10 multiple choice problems on your test, with 4 choices each. What is the probability that you will get a minimum passing grade?
Minimum passing grade would be a 60% and above

Answer by stanbon(75887)

There are 10 multiple choice problems on your test, with 4 choices each. What is the probability that you will get a minimum passing grade?
Minimum passing grade would be a 60% and above
--------------------
>=60% means >=6 correct problems.
------------------------------
Binomial Problem with n = 10 and p = 1/4
---
P(x >= 6) = 1 - P(0<= x<=5) = 1-binomcdf(10,1/4,5)
----
= 0.019728
===============
Cheers,
Stan H.
===============